how would I find the Slope Intercept From of this equation: 4x-5y=15?

Answer by ewatrrr(24785)


Hi,
Solve for y to put into the slope intercept form of an equation for a line
y = mx + b where m is the slope and b the y intercept
.
4x-5y=15
4x - 15 = 5y
(4/5)x - 3 = y
y = (4/5)x - 3
slope of this line = (4/5)